# Bipartite Discrete Time Crystals on Decorated Lattices

### v1

@@ Abstract
As seen from the spectral density $|S_{\pi/\tau}^{B}|^{2}$ of the subharmonic response in the thermodynamic limit of a Lieb lattice ($z=4$) and decorated 5-regular graph ($z=5$) shown in Fig. 4(c-d), the recurring appearance and disappearance of DTC order as a function of $\theta_{J}$ is a general feature of decorated lattice structures. In particular, nodes of vanishing DTC order occur at values $\theta_{J}=-2n\pi/z$, with $n\in\mathbb{N}$ and $z$ the connectivity of the $B$-lattice. This is intuitively explained by the fact that the effective rotation experienced by a $z$-connected qubit in the $B$-lattice upon the application of $z$ $R_{ZZ}$ gates is then $2n\pi$, reintroducing the instability of a non-interacting model. By the same reasoning, the order parameter in the $A$-sublattice of any decorated lattice reaches a node at $\theta_{J}=-\pi$, as in Fig. 4(a). However, while nodes of the $B$-lattice occur also in the phase diagram of the $A$-lattice and thus destabilize DTC order in the entire system, the $A$-lattice node at $-\pi$ results in a regime of bipartite DTC order: qubits in the $B$-lattice [Fig. 4(b)] exhibit a robust subharmonic response in the limit $\theta_{J}\to-\pi$, whereas the $A$-lattice [Fig. 4(a)] acquires a quasi-crystalline beating similar to the dynamics in Fig. 3. This is shown by the blue-shaded amplitude of the largest non-subharmonic peak in the spectral density $|S_{\omega}^{A}|^{2}$ in Fig. 4(a). Finally, we note that bipartite DTCs do not appear in decorated square lattices with an even coordination number $z=4$ in the $B$ lattice. This follows from the above argument and is visible from the $B$-lattice order parameter in the Lieb lattice shown in Fig. 4(c), where the coincidence of DTC nodes in both sublattices at $\theta_{J}=-\pi$ inhibits the revival of bipartite DTC order witnessed in decorated hexagonal and 5-regular graphs.
@@
Conclusion and outlook. —  Through tensor network simulations in the thermodynamic limit, we have studied discrete time crystals in periodically driven Ising models on decorated lattices. We have established that prethermal DTC order decays in a non-perturbative fashion from small deviations from the ideal $\pi$-periodic drive, and have found that the sensitivity to a longitudinal field, observed on a finite decorated hexagonal lattice in Ref. [37], manifests itself in only one of the two sublattices comprising a generic decorated graph structure. The latter observation implies the existence of a state of bipartite time-crystalline order, in which an extensive part of a bipartite system resisting thermalization displays subharmonic dynamics robust to generic local perturbations. Finally, we have constructed a prethermal DTC phase diagram for the specific cases of a decorated hexagonal lattice, Lieb lattice and decorated 5-regular lattice, uncovering a decay and resurrection of (bipartite) DTC order at interaction strengths determined by the topology of the graph.
@@
The bipartite discrete time crystals we have introduced bear a resemblance to the notion of boundary time crystals (BTC) [48, 49, 50] and topological Floquet phases of edge states [51]. In contrast to these phenomena, bipartite DTCs as introduced in this work are distinguished by their extensive scaling with system size. As such, we anticipate experimental realizations of this phenomenon not only on current digital quantum processors [52, 53, 51, 54, 55, 56], but also on atomic systems with tunable lattice structures [57, 58, 59] and condensed matter platforms with non-trivial topologies, such as spin impurities in diamond [60, 47]. Finally, additional research is required to explore whether the phenomenology presented here can be extended to the breaking of a continuous time symmetry in driven-dissipative systems [61, 62, 63, 64] with a bipartite geometry.
Acknowledgements.
@@
Appendix A Convergence of BP-iTNS results
@@
To validate our BP-iTNS results obtained for the unit cell of an infinite lattice, we provide in Figure 5(a) a scaling analysis of the result of Fig. 2 for finite systems composed of $L\times L$ decorated hexagons for $L=\{1,2,3,4\}$, corresponding to $N=\{12,35,68,111\}$ qubits. As proven in more detail in Ref. [33], this illustrates that the BP approximation converges rapidly with system size in both the time-crystalline and thermalizing regimes. The inset shows a repetition of Fig. 3(c) but measured in the bulk of a $111$-site system composed of $4\times 4$ decorated hexagons, illustrating the survival of bipartite DTC order in finite systems. In infinite lattices represented by the $K_{2,z}$ unit cell, we have confirmed that bipartite DTC order is robust w.r.t. symmetry-broken initial states of the form $\bigotimes_{i}^{z+2}|\sigma_{i}\rangle$, with $\sigma_{i}=\uparrow\downarrow$ the $Z$-polarized state of a qubit in the unit cell.
@@
As dicussed in the main text and shown in Fig. 2(e), the lifetime of the prethermal DTC scales exponentially with $1/|\theta_{x}-\pi|$. To further illustrate the convergence of our results during the prethermal stage, we show in Fig. 5(b) the mean magnetization in the $N$-qubit system averaged over several numbers $t/\tau$ of Floquet cycles,
$\overline{Z}=\frac{1}{n_{\tau}N}\sum_{n=1}^{t/\tau}\sum_{j=1}^{N}\langle\psi(n\tau)|(-1)^{n}Z_{j}|\psi(n\tau)\rangle.$ (5)
@@
While the magnetization converges rapidly in the DTC regime $\theta_{x}\to\pi$, the thermalizing regime is characterized by a slow decay, consistent with the eventual disappearance of DTC order at long times. The field strength $\theta_{x}^{*}$ of the crossover between these regimes, indicated in Fig. 2(d), is inferred from the convergence in time of the finite entanglement scaling. As shown in Fig. 5(c), the slope of the extrapolation reaches a maximum in the intermediary regime, which converges to $\theta_{x}^{*}\approx 0.73$ with increasing number of Floquet cycles.
Figure 5: Scaling analysis of BP-iTNS results. (a) Scaling of the time-averaged magnetization over 100 Floquet cycles with the number of unit cells in a finite decorated hexagonal lattice evaluated at $\chi=80$. The inset shows the same result as Fig. 3(c) evaluated in the bulk of a $111$-site system composed of $4\times 4$ decorated hexagons. (b) Convergence of the magnetization in the BP-iTNS with the number of cycles at $\chi=160$. (d) Evolution of the slope in the $\chi\to\infty$ extrapolation shown in Fig. 2.
Appendix B State vector simulation
@@
For small systems, the dynamics generated by the Floquet unitary can be evaluated exactly by the direct application of one- and two-qubit gates on the state vector expressed in the computational basis of the $2^{N}$-dimensional Hilbert space. This allows us to verify our conclusions obtained in the BP-iTNS approach with small graphs exhibiting a similar topology. To validate the existence of bipartite DTCs, we simulate the Floquet dynamics considered in Fig. 3 of the main text on complete $z$-regular graphs of a minimal size for $z=\{3,4,5\}$, with decorated edges. Shown in the top row are the magnetization dynamics on any of the equivalent qubits in the undecorated graphs comprising only the red nodes. For Floquet parameters $(\theta_{J},\theta_{x},\theta_{z})=(0.5\pi,0.85\pi,0.4\pi)$, the magnetization lacks robust subharmonic order for graphs of any degree, and instead experiences a beating due to the presence of both a transverse and longitudinal field. Decorating the edges with additional (orange) qubits gives rise to the bipartite graphs shown in the top row of Fig. 6. For $z=3$ (a) and $z=4$ (b), these graphs have the same degrees of connectivity as the iTNS unit cells shown in Fig. 1. As in the main text, the resulting $A$- and $B$-sublattices now exhibit qualitatively different dynamics. Whereas the dynamics of the magnetization on the decorated qubits is unstable to any longitudinal field perturbation, the subharmonic response on $z$-connected qubits is stabilized for $z=3$ [Fig. 6(a)] and $z=5$ [Fig. 6(c)]. By contrast, time-crystalline order in the magnetization of a $z=4$ bipartite graph [Fig. 6(b)] appears uniformly across both sublattices, as it does in the BP-iTNS approximation of an infinite Lieb lattice.
@@
Figure 6: Bipartite rigidity in decorated regular graphs. Exact dynamics of the magnetization $\langle Z\rangle$ generated by the Floquet unitary (3) with rotation angles $(\theta_{x},\theta_{z})=(0.85\pi,0.4\pi)$ as in Fig. 3(c), on a $3$-regular (a), $4$-regular (b) and $5$-regular graph. The top row shows the dynamics for the undecorated graphs (i.e., in absence of the orange nodes); the bottom row shows the dynamics on the $A$ and $B$-qubits of the bipartite decorated graphs. The interaction strength $\theta_{J}$ was taken to be $\pi/2$ (a,c) and $\pi/3$ (b), within the DTC regime of the respective decorated graphs.

### v2

@@ Abstract
For the decorated hexagonal lattice in Fig. 4(a-b), the stabilizing effect of interactions stagnates around $\theta_{J}\approx-\pi/3$. Beyond this point, the range of perturbations $|\theta_{x}-\pi|$ to which the time crystal is robust shrinks until the eventual disappearance of DTC order in the node at $\theta_{J}=-2\pi/3$, followed by a revival at larger interaction strengths. This disappearance and recurrence of DTC order as a function of $\theta_{J}$ is a general feature of decorated lattice structures, as seen from the $B$-sublattice order parameter $|S_{\pi/\tau}^{B}|^{2}$ in the thermodynamic limit of a Lieb lattice ($z=4$) and decorated 5-regular graph ($z=5$) shown in Fig. 4(c-d). In particular, nodes of vanishing DTC order occur at values $\theta_{J}=-2n\pi/z$, with $n\in\mathbb{N}$ and $z$ the connectivity of the $B$-lattice. This property is readily explained by the fact that the $R_{ZZ}^{jk}(\theta_{J})$ interaction terms in Eq. (3) effectively rotate each of the involved qubits by an angle $\theta_{J}$ around the $Z$-axis. Hence, the total rotation of a $z$-connected qubit in the $B$-lattice interacting with its $z$ neighbors in a single Floquet cycle is $z\theta_{J}$. At $\theta_{J}=-2n\pi/z$ this reduces to $-2n\pi$, reintroducing the instability of a non-interacting model at $\theta_{J}=0$. Following the same reasoning, qubits in the $A$-sublattice of any decorated lattice have two nearest neighbors and thus experience an effective rotation of $2\theta_{J}$. As seen in Fig. 4(a), this results in the disappearance of DTC order in the $A$-sublattice at interaction strengths $\theta_{J}=-n\pi$.
@@
However, while nodes of the $B$-lattice occur also in the phase diagram of the $A$-lattice and thus destabilize DTC order in the entire system, the $A$-lattice node at $\theta_{J}=-\pi$ results in a regime of bipartite DTC order: qubits in the $B$-lattice [Fig. 4(b)] exhibit a robust subharmonic response in the limit $\theta_{J}\to-\pi$, whereas the $A$-lattice [Fig. 4(a)] acquires a quasi-crystalline beating similar to the dynamics in Fig. 3. This is shown by the blue-shaded amplitude of the largest non-subharmonic peak in the spectral density $|S_{\omega}^{A}|^{2}$ in Fig. 4(a). Finally, we note that regimes of bipartite DTC order are not observed in Lieb lattices with even coordination number $z=4$. This follows from the counting argument above, according to which the $B$-lattice exhibits a DTC node at $\theta_{J}=-n\pi/2$, shown in Fig. 4(c). The coincidence of the DTC node $\theta_{J}=-\pi$ with that of the $A$-lattice prevents the emergence of a bipartite DTC phase witnessed in decorated hexagonal and 5-regular graphs.
@@
Conclusion. —  Through tensor network simulations in the thermodynamic limit, we have studied discrete time crystals in periodically driven Ising models on decorated lattices. We have established the existence of a prethermal phase of time-crystalline order which is exponentially long-lived and decays in a nonperturbative fashion from small deviations from the ideal $\pi$-periodic drive. Additionally, we have found that the sensitivity to a longitudinal field, observed on a finite decorated hexagonal lattice in Ref. [49], manifests itself in only one of the two sublattices comprising a generic decorated graph structure. The latter observation implies the existence of a state of bipartite time-crystalline order, in which an extensive part of a bipartite system resisting thermalization displays subharmonic dynamics robust to generic local perturbations. Finally, we have constructed a prethermal DTC phase diagram for the specific cases of a decorated hexagonal lattice, Lieb lattice and decorated 5-regular graph, uncovering a decay and resurrection of both uniform and bipartite prethermal DTC order at interaction strengths determined by the topology of the graph.
@@
The bipartite discrete time crystals we have introduced bear a resemblance to the notion of boundary time crystals (BTC) [75, 76, 77] and topological Floquet phases of edge states [78, 79, 50]. In contrast to these phenomena, bipartite DTCs as introduced in this work are distinguished by their extensive scaling with system size. As such, we anticipate experimental realizations of this phenomenon not only on current digital quantum processors [80, 81, 82, 78, 83, 84, 48], but also on analog quantum simulators with periodically driven Ising interactions on non-trivial topologies. Possible candidates for such studies include atomic systems with tunable lattice structures [85, 86, 87] and condensed matter platforms such as spin impurities in diamond [22, 73]. However, further theoretical research is required to elucidate the physical mechanisms underpinning bipartite time-crystallinity, as well as to investigate whether the phenomenology presented here can be extended to the breaking of a continuous time symmetry in driven-dissipative systems [88, 89, 90, 91] with a bipartite geometry.
@@
Supplemental material
to
Nonperturbative decay of bipartite discrete time crystals
Lennart Fernandes, Joseph Tindall, and Dries Sels
I Convergence of BP-iTNS results
@@
To validate our BP-iTNS results obtained for the unit cell of an infinite lattice, we provide in Figure S1(a) a scaling analysis of the result of Fig. 2(d) in the main text for finite systems composed of $L\times L$ decorated hexagons for $L=\{1,2,3,4\}$, corresponding to $N=\{12,35,68,111\}$ qubits. As proven in more detail in Ref. [61], this illustrates that the BP approximation converges rapidly with system size in both the time-crystalline and thermalizing regimes. The inset shows a repetition of Fig. 3(c) in the main text, but measured in the bulk of a $111$-site system composed of $4\times 4$ decorated hexagons, illustrating the survival of bipartite DTC order in finite systems. In infinite lattices represented by the $K_{2,z}$ unit cell, we have confirmed that bipartite DTC order is robust w.r.t. symmetry-broken initial states of the form $\bigotimes_{i=1}^{z+2}|\sigma_{i}\rangle$, with $\sigma_{i}=\uparrow\downarrow$ a $Z$-polarized state of a qubit in the unit cell. We note that the employed BP-iTNS approach intrinsically limits our results in the thermodynamic limit to systems with spatial translational invariance.
@@
As dicussed in the main text, the lifetime of the prethermal DTC scales exponentially with $1/|\theta_{x}-\pi|$. To illustrate the convergence of our results within a long prethermal stage, we show in Fig. S1(b) the mean magnetization in the $N$-qubit system averaged over several numbers $t/\tau$ of Floquet cycles,
$\overline{Z}=\frac{\tau}{tN}\sum_{n=1}^{t/\tau}\sum_{j=1}^{N}\langle\psi(n\tau)|(-1)^{n}Z_{j}|\psi(n\tau)\rangle.$ (S1)
@@
The thermalizing regime is characterized by a persistent decay of magnetization consistent with the eventual disappearance of order at long times. By contrast, in the DTC regime $\theta_{x}\to\pi$, $\overline{Z}$ shows no significant decay, highlighting the persistence of DTC order over long time scales despite perturbations. The field strength $\theta_{x}^{*}$ of the crossover between these regimes, indicated in Fig. 2(d-e) in the main text, is inferred from the convergence in time of the finite entanglement scaling relation Eq. (5) in the main text. The validity of this linear scaling in $1/\ln(\chi)$, illustrated by the inset of Fig. 2(d) of the main text, is shown in Fig. S2(a) for a wider range of $\theta_{x}$ in both the thermalizing and DTC regimes. As shown in Fig. S2(b), the extrapolation slope $\zeta(\theta_{x})$ reaches a maximum in the intermediary regime, which converges to $\theta_{x}^{*}\approx 0.73$ with increasing number of Floquet cycles.
II State vector simulation
@@
For small systems, the dynamics generated by the Floquet unitary can be evaluated exactly by the direct application of one- and two-qubit gates on the state vector expressed in the computational basis of the $2^{N}$-dimensional Hilbert space. This allows us to verify our conclusions obtained in the BP-iTNS approach with small graphs exhibiting a similar topology. To validate the existence of bipartite DTCs, we simulate the Floquet dynamics considered in Fig. 3 of the main text on complete $z$-regular graphs of a minimal size for $z=\{3,4,5\}$, with decorated edges. Shown in the top row are the magnetization dynamics on any of the equivalent qubits in the undecorated graphs comprising only the red nodes. For Floquet parameters $(\theta_{J},\theta_{x},\theta_{z})=(0.5\pi,0.85\pi,0.4\pi)$, the magnetization lacks robust subharmonic order for graphs of any degree, and instead experiences a beating due to the presence of both a transverse and longitudinal field. Decorating the edges with additional (orange) qubits gives rise to the bipartite graphs shown in the top row of Fig. S3. For $z=3$ (a) and $z=4$ (b), these graphs have the same degrees of connectivity as the iTNS unit cells shown in Fig. 1(c-d). As in the main text, the resulting $A$- and $B$-sublattices now exhibit qualitatively different dynamics. Whereas the dynamics of the magnetization on the decorated qubits is unstable to any longitudinal field perturbation, the subharmonic response on $z$-connected qubits is stabilized for $z=3$ [Fig. S3(a)] and $z=5$ [Fig. S3(c)]. By contrast, time-crystalline order in the magnetization of a $z=4$ bipartite graph [Fig. S3(b)] appears uniformly across both sublattices, as it does in the BP-iTNS approximation of an infinite Lieb lattice.
Figure S1: Scaling analysis of BP-iTNS results. (a) Scaling of the time-averaged magnetization over 100 Floquet cycles with the number of unit cells in a finite decorated hexagonal lattice evaluated at $\chi=80$. The inset shows the same result as Fig. 3(c) evaluated in the bulk of a $111$-site system composed of $4\times 4$ decorated hexagons. (b) Convergence of the magnetization in the BP-iTNS with the number of cycles at $\chi=320$.
Figure S2: Finite entanglement scaling. (a) Linear scaling of the time-averaged magnetization during $t/\tau=100$ Floquet cycles with $1/\ln(\chi)$, as in the inset of Fig. 2(d), for $\theta_{x}/\pi=\{0.6:0.05:1.0\}$ (light to dark). (b) Time evolution of the extrapolation slope $\zeta(\theta_{x})$ as in Fig. 2(e), emphasizing the convergence of the maximum to $\theta_{x}^{*}$.
@@
Figure S3: Bipartite rigidity in decorated regular graphs. Exact dynamics of the magnetization $\langle Z\rangle$ generated by the Floquet unitary with rotation angles $(\theta_{x},\theta_{z})=(0.85\pi,0.4\pi)$ as in Fig. 3(c), on a $3$-regular (a), $4$-regular (b) and $5$-regular graph. The top row shows the dynamics for the undecorated graphs (i.e., in absence of the orange nodes); the bottom row shows the dynamics on the $A$ and $B$-qubits of the bipartite decorated graphs. The interaction strength $\theta_{J}$ was taken to be $\pi/2$ (a,c) and $\pi/3$ (b), within the DTC regime of the respective decorated graphs.